How many milliliters of water must be added to 10 milliliters of a 30%-sulfuric acid solution to make a 15%-sulfuric acid solution?
Answer by mananth(16946)

30% sulphuric acid = 70% water
15% sulphuric acid = 85% water
Let quantity of water be x ml
concentration equation
x+0.7*10 = 0.85(x+10)
x+7=0.85x+8.5
0.15x=1.5
x= 1.5/0.15
x=10 ml